# UAV Path Planning for Object Observation with Quality Constraints: A Dynamic Programming Approach

### v1

@@ II-C Integer Linear Programming representation
Based on the discrete observation points discussed in Section II-B, an Integer Linear Programming (ILP) approach is proposed. The approach is based on the concept of dividing the map into independent zones, each comprising an object’s effective observing area and observation points capable of covering it. The starting point is considered an independent zone that contains only a single point. The subsequent step involves selecting a single point from each zone to form a route that passes through all observation points and the starting point. The formulated ILP approach is based on the Traveling Salesman Problem (TSP) formulation and is illustrated in Figure. 3.
@@
$\sum_{j\notin\{1,N+1\}}\sum_{p_{2}}(X_{1,j,p_{1},p_{2}}+X_{N+1,j,p_{1},p_{2}})$ $=\sum_{j\notin\{1,N+1\}}\sum_{p_{2}}(X_{j,1,p_{2},p_{1}}+X_{j,N+1,p_{2},p_{1}})$ $\forall p_{1}\in zone1$ (9)
@@
$\sum_{i}\sum_{p_{1}}q_{i,p_{1}}\sum_{j}\sum_{p_{2}}X_{i,j,p_{1},p_{2}}$ $*\displaystyle\geq q*$ (11)
$X_{ijp_{1}p_{2}}\in\left\{0,1\right\},u_{i}$ $\geq 0,u_{i}\in R$ (12)
Figure 3: ILP formulation
@@ III Methodology
III-A Dynamic programming adjustment
A dynamic programming algorithm, the core part of our work, is presented at length in this section. More specifically, it processes from the first object to the last one in order by continuously updating the shortest distance from the starting point to the current point, as the relaxation does in Dijkstra’s algorithm [38]. The quality constraint necessitates recording the observation quality at each point along the path, which leads to the design of a multidimensional table. Furthermore, the shortest path to $o_{i}$ can be exactly a straight line from that of $o_{j}$ if there is an overlap of the observation areas of objects $o_{j+1}$ to $o_{i}$ so that one observation point may be sufficient for the UAV to cover all of them (Figure. 4).
Definition 2.
Let $dp[i,p_{i}]$ be the set containing pairs of length and observation quality $(d,q)$, which implies one path to observation points $p_{i}$ of the $i^{th}$ objects. Formally, $dp[i,p_{i}]=\{(d_{1},q_{1}),(d_{2},q_{2}),...,(d_{z},q_{z})\}$.
Note that $i$ here is the index of the object in the visiting order. If the order is $0,1,3,4,2,0$ in which 00 refers to the starting point, the second object to observe is $o_{3}$.
Definition 3.
@@ III-A Dynamic programming adjustment
Lemma 2.
The number of observation points for $n$ objects is $O(\frac{n^{3}{\epsilon}+n^{2}}{\epsilon^{3}})$.
Proof.
Provided by Theorem 4.4 in [36], the number of observation points for $n$ objects is $O(\frac{n^{2}}{\epsilon^{3}})$. By using the discretization method of $\delta$, the discretization interval is $\frac{{\epsilon}D}{n}$. In this case, there are at most $\frac{d_{max}n}{{\epsilon}D}$ points in the distance domain and at most $\frac{d_{max}{\theta}n}{{\epsilon}D}$ points in the angle domain, the multiple of the which is the maximum number of observation points for one object. Therefore, the number of all the observation points $|\mathcal{P}|$ is:
$O\left(\frac{n^{2}}{\epsilon^{3}}\right)+O\left(\frac{nd_{max}}{\epsilon{D}}\cdot\frac{nd_{max}{\theta}}{\epsilon{D}}\cdot n\right)$ (18)
$=$ $O\left(\frac{n^{2}}{\epsilon^{3}}+\frac{n^{3}}{\epsilon^{2}}\right)$ (19)
Theorem 1.
@@
$O\left(\log(\frac{n(D+2d_{max})}{\delta})\right)$ (20)
$=$ $O\left(\log(\frac{n^{2}(D+2d_{max})}{{\epsilon}D})\right)$ (21)
$=$ $O\left(\log(\frac{n^{2}}{\epsilon})\right)$ (22)
@@
$O\left(n\cdot\frac{|\mathcal{P}|}{n}\cdot n\cdot\frac{|\mathcal{P}|}{n}\cdot\log(\frac{n^{2}}{\epsilon})\right)$ (23)
$=$ $O\left((\frac{n^{2}+n^{3}\epsilon}{\epsilon^{3}})^{2}\cdot\log(\frac{n^{2}}{\epsilon})\right)$ (24)
III-B Determination of observing order
@@ III-B1 Random Selection (RS)
III-B2 Nearest Point First (NPF)
The second method is setting off from the starting point and always choosing the nearest observation point that belongs to an object not visited yet. The process is continuously repeated until all objects are observed. The UAV then directly returns to the starting point.
@@ III-B2 Nearest Point First (NPF)
5:     choose a nearest point $p_{next}$ from $\mathcal{P}$
@@ III-B3 Generalized TSP (GTSP)
III-B4 TSP over Objects (TSPO)
@@
III-B5 TSP on Lower Bound (LBTSP)
Given the proposed lower bound in Section III-C, there is a weighted graph essentially including all the objects and the starting point $P_{0}$. Even though the graph may not be subject to the Euclidean metrics, we can still obtain a visiting order by conducting the TSP algorithm on it.
III-C Lower bound
To calculate the approximation ratio of the algorithms in experiments, a lower bound of the problem is designed and is described in this part.
We consider the observation points of each object to be one cluster and $P_{0}$ as an independent cluster, similar to the design of ILP in Section II-C. Then we consider each cluster to be a single vertex after solving the minimum distance between each of them. If there is an overlapping area between two clusters, the minimum distance between them is 0. This distance is used as the weight of the edges among vertices, hence a graph is formed. The lower bound is exactly the cost of the minimum spanning tree of the graph.
@@ III-C Lower bound
IV Experiments
In this section, we evaluate the performance of the proposed algorithms via carrying out a series of simulated experiments. The experiments contain two parts, the first of which focuses on the numerical results of synthetic cases while the second is conducted in the Airsim simulator[40] to test the performance in the realistic virtual environment.
IV-A Experiments on synthetic cases
@@
IV-A1 Brute force experiments
@@
IV-A2 Different range and angle experiments
@@
IV-B Experiments in the simulator
Besides the tests on synthetic data, we carried out our algorithms in the Airsim simulator (see Figure. 5), which is an open-source simulator developed by Microsoft for training and testing autonomous vehicles and drones in realistic virtual environments. We create 50 cylinders as objects in the 3dsmax software with radius of 1m and height of 2m. After that, we paste on the cylinders high definition English advertisement images downloaded from the Internet, which are then imported into the simulator. The map size is set to 150m and 30 cases are generated for each $n$ in $\left\{5,10,15,20\right\}$. For each case, $n$ objects are randomly chosen from the 50 ones and $q^{*}$ is set to $\left\{0.3,0.5,0.7,0.9\right\}$. The drone will fly along the paths produced by different order determination algorithms and dynamic programming algorithm. Additionally, the initial paths are included provided by RS, NPF and GTSP without adjustment. We also add a path that yields the highest observation quality from all the objects. The path is solved by using the algorithm [32] on the observation points that produces maximum observation quality of each object. During the flight, the UAV takes photos of the objects which are sent to the CnOCR API in Python [41]. The observation quality can thus be reflected from the accuracy of the recognized words in the photos.
@@ IV-B Experiments in the simulator
To calculate the recognition accuracy in one case, we solve the longest common subsequence between the recognition result and the label of the object as the correctly recognized string. Suppose that the labels of the objects in one case are =$\mathcal{L}=\left\{l_{1},l_{2},...,l_{m}\right\}$, the correctly recognized strings $\mathcal{C}=\left\{c_{1},c_{2},...,c_{m}\right\}$. We define the accuracy as the sum of all correctly recognized letters over all the letters in the labels, $\frac{\sum_{i}{|c_{i}|}}{\sum_{i}{|l_{i}|}}$. Figure. 6 shows the photos taken during the flight and the according OCR results.
Refer to caption (a) Refer to caption (b) Refer to caption (c)
The first section of experiments consider that the observation distance and angle are equal for all objects. In the simulation trails, however, owing to the different size of the images and words, the observing distance and observing angle are calculated separately for each object. We stipulate that the UAV should recognize at least $30\%$ correct letters for each object from some observation point. Based on this, $d_{max}$ and $d_{min}$ are searched via a binary search algorithm, after placing the UAV right on the front of the object. The observing angle $\theta$ is then obtained from a second binary search algorithm that guides the UAV to fly on the arc with the radius of $d_{max}$. Table I contains a proportion of the objects’ information.
TABLE I: Labels, maximum and minimum observation distance and maximum observation angles of some objects
@@
IV-C Numerical results
In this part, the statistic results are presented in the following figures and tables. We first present the results of the Brute force experiments, which are the comparison of our order determination algorithms with the optimal order. The results of different range and angle experiments are then demonstrated before presenting the running time. Finally, we show the simulation results.
IV-C1 Comparison with brute force
@@
IV-C2 Different $n$ and range
We fix the observation range to 4, 8 and 12 and plot Figure. 7, in which the effect of $n$ is illustrated. After that, $n$ is fixed and the effect of observation range is shown in Figure. 8.
As when $d_{max}=4$, the overall approximation ratio has a declining tendency with 1.61 when $n$ is 5 and 1.31 when $n$ is 30 for the GTSP method. Other order determination methods produce similar results and the relative superiority is shown to be the same as that in the brute force section. NPF achieves 1.71 approximation ratio when $n$ is 5 which gradually declines to 1.58 when $n$ is 30. The results bear some differences for larger observation range. When the range is 8 and 12, the approximation ratios both undergo a decline when $n$ is smaller than 15 and then experience a rebound. Concretely, for cases where the range equals 8, The GTSP methods achieves the approximation ratio of 1.69 when $n$ is 5, which gradually descends to 1.45 when $n$ is 15. The ratio then rebound to 1.49. When $n$ is 12, the approximation ratio ranges from 1.51 to 1.98, slightly higher than the overall situation in $n=8$. The performance of NPF method seems to be susceptible to the growing $n$ but the other method remains relatively stable. In another word, our algorithms are robust to the growing number of objects.
Refer to caption (a) Refer to caption (b) Refer to caption (c)
@@ IV-C2 Different $n$ and range
We then analyse the effect of observation range. The number of objects are selected to be respectively 10, 20 and 30. It is apparent to notice the growing trend in approximation ratio with the rise in the observation range from 4 to 10 for all $n$. The ratio then remains stable when the range further increases to 12. When $n$ is 10, GTSP method achieves the approximation ratio ranging from 1.41 to 1.59, roughly 0.1 or 0.2 smaller than the other methods. The results bear a resemblance for large n, in which the ratio increases from 1.32 to 1.52 when $n$ is 20 and 1.32 to 1.6 when $n$ is 30. As a brief summary, there is a modest increase in approximation ratio along with the observation range and number of objects, in that the algorithms probably have difficulty in solving complex cases.
Refer to caption (a) Refer to caption (b) Refer to caption (c)
@@
IV-C3 Effectiveness of dynamic programming algorithm
@@
We have counted the percentage of cases in which RS, NPF and GTSP methods can achieve the quality requirement without the assistance from our dynamic programming algorithm, the effectiveness and the necessity of which is to be explored. TSPO and LBTSP methods are omitted since they only yield the visiting order instead of feasible paths. For those cases in which the order determination methods have already been capable of achieving the quality requirement, we calculate the percentage by which the dynamic programming algorithm is able to reduce the path length.
@@
The number of objects is selected to be 5, 15 and 25. As the quality requirement grows, the general tendency is extremely in accordance with the expectation. Concretely, the satisfying percent of all the three methods remains high when $q^{*}\leq 0.5$, with GTSP order performing slightly poorly than the others. For stricter $q^{*}$, however, the percentage drops rapidly to 20% and finally to 0, which is given by NPF method when $q^{*}$ is 0.9. Based on the outcomes above that the order determination algorithms can barely find satisfying paths for large $q^{*}$, it is necessary to use dynamic programming to improve the path in order to fulfill the quality requirement.
After rendering the path to achieve enough observation quality, our adjustment algorithm can further reduce the path length to some extent. Regarding the reduction in initial paths, the reduced percent of the length increases with $n$ increasing, from around 3. 5%, to 5. 5% and further to 6% when $n$ is 25. From another perspective, the reduction rates keep steady in all of the three methods when $q^{*}$ is less than 0.6, with the reduction in RS close to that in NPF and more significant than that in GTSP. Nevertheless, the reduction slows down distinctly as $q^{*}$ increases additionally, given the available space is becoming smaller for the UAV. There is an exception that the reduced percent of GTSP method remains steady regardless of the alternation in both $q^{*}$ and $n$. The results on max quality paths bear more significant improvement with around 16% when $n$ is 5. The value decreases slightly to around 15% and 12% when $n$ grows to 15 and 25. In seeking a comparatively good observation quality, there is a worthwhile trade-off on the quality and the path length. For instance, the algorithm can seek the path that has 90% observation quality but an evident reduction in the length cost.
IV-C4 Simulation results

### v2

@@ 3.3 Integer Linear Programming representation
Based on the discrete observation points discussed in Section 3.2, an Integer Linear Programming (ILP) approach is proposed. The approach is based on the concept of dividing the map into independent zones, each comprising an object’s effective observing area and observation points capable of covering it. The starting point is considered an independent zone that contains only a single point. The subsequent step involves selecting a single point from each zone to form a route that passes through all observation points and the starting point. The formulated ILP approach is based on the Traveling Salesman Problem (TSP) formulation and is illustrated in Equation (5) to Equation (12).
@@
$\sum_{j\notin\{1,N+1\}}\sum_{p_{2}}(X_{1,j,p_{1},p_{2}}+X_{N+1,j,p_{1},p_{2}})$
$\quad=\sum_{j\notin\{1,N+1\}}\sum_{p_{2}}(X_{j,1,p_{2},p_{1}}+X_{j,N+1,p_{2},p_{1}})$ $\forall p_{1}\in zone1$ (9)
@@
$\sum_{i}\sum_{p_{1}}q_{i,p_{1}}\sum_{j}\sum_{p_{2}}X_{i,j,p_{1},p_{2}}\geq q*$ (11)
$X_{i,j,p_{1},p_{2}}\in\left\{0,1\right\},u_{i}\geq 0,u_{i}\in R$ (12)
@@ 4 Methodology
4.1 Dynamic programming adjustment
A dynamic programming algorithm, the core part of our work, is presented at length in this section. More specifically, it processes from the first object to the last one in order by continuously updating the shortest distance from the starting point to the current point, as the relaxation does in Dijkstra’s algorithm [38]. The quality constraint necessitates recording the observation quality at each point along the path, which leads to the design of a multidimensional table. Furthermore, the shortest path to $o_{i}$ can be exactly a straight line from that of $o_{j}$ if there is an overlap of the observation areas of objects $o_{j+1}$ to $o_{i}$ so that one observation point may be sufficient for the UAV to cover all of them (Figure 3).
Definition 2
Let $dp[i,p_{i}]$ be the set containing pairs of the path length and observation quality $(d,q)$, which implies one path to observation points $p_{i}$ of the $i^{th}$ objects. Formally, $dp[i,p_{i}]=\{(d_{1},q_{1}),(d_{2},q_{2}),...,(d_{z},q_{z})\}$.
Note that $i$ here is the index of the object in the visiting order. If the order is $[0,1,3,4,2,0]$ in which 00 refers to the starting point, the second object to observe is $o_{3}$.
Definition 3
@@ 4.1 Dynamic programming adjustment
Lemma 2
The number of observation points for $n$ objects is $O\left(\frac{n^{3}{\epsilon}+n^{2}}{\epsilon^{3}}\right)$.
Proof 2
Provided by Theorem 4.4 in [36], the number of observation points for $n$ objects is $O\left(\frac{n^{2}}{\epsilon^{3}}\right)$. By using the discretization method of $\delta$, the discretization interval is $\frac{{\epsilon}D}{n}$. In this case, there are at most $\frac{d_{max}n}{{\epsilon}D}$ points in the distance domain and at most $\frac{d_{max}{\theta}n}{{\epsilon}D}$ points in the angle domain, the multiple of the which is the maximum number of observation points for one object. Therefore, the number of all the observation points $|\mathcal{P}|$ is:
$O\left(\frac{n^{2}}{\epsilon^{3}}\right)+O\left(\frac{nd_{max}}{\epsilon{D}}\cdot\frac{nd_{max}{\theta}}{\epsilon{D}}\cdot n\right)=O\left(\frac{n^{2}}{\epsilon^{3}}+\frac{n^{3}}{\epsilon^{2}}\right)$ (16)
Theorem 1
@@
$O\left(\log(\frac{n(D+2d_{max})}{\delta})\right)=$ $O\left(\log(\frac{n^{2}(D+2d_{max})}{{\epsilon}D})\right)=$ $O\left(\log(\frac{n^{2}}{\epsilon})\right)$ (17)
@@
$O\left(n\cdot\frac{|\mathcal{P}|}{n}\cdot n\cdot\frac{|\mathcal{P}|}{n}\cdot\log(\frac{n^{2}}{\epsilon})\right)=O\left((\frac{n^{2}+n^{3}\epsilon}{\epsilon^{3}})^{2}\cdot\log(\frac{n^{2}}{\epsilon})\right)$ (18)
4.2 Determination of observing order
@@ 4.2.1 Random Selection (RS)
4.2.2 Nearest Point First (NPF)
The second method (shown in Algorithm 2) is setting off from the starting point and always choosing the nearest observation point that belongs to an object not visited yet. The process is continuously repeated until all objects are observed. The UAV then directly returns to the starting point.
@@ 4.2.2 Nearest Point First (NPF)
5:     choose the nearest point $p_{next}$ from $\mathcal{P}$
@@ 4.2.3 Generalized TSP (GTSP)
4.2.4 TSP over Objects (TSPO)
@@
4.2.5 TSP on Lower Bound (LBTSP)
Given the proposed lower bound in Section 4.3, there is a weighted graph essentially including all the objects and the starting point $P_{0}$. Even though the graph may not be subject to the Euclidean metrics, we can still obtain a visiting order by conducting the algorithm from [32] on it.
4.3 Lower bound
To calculate the approximation ratio of the algorithms in experiments, a lower bound of the problem is designed and is described in this part.(shown in Algorithm 3)
We consider the observation points of each object to be one cluster and $P_{0}$ as an independent cluster, similar to the design of ILP in Section 3.3. Each cluster is then viewed as a single vertex after solving the minimum distance between each of them. If there is an overlapping area between two clusters, the minimum distance between them is 0. This distance is used as the weight of the edges among vertices, hence a graph is formed. The lower bound is exactly the cost of the minimum spanning tree of the graph.
@@ 4.3 Lower bound
1:Object set $\mathcal{O}$, Observation point set $\mathcal{P}$, and starting point $P_{0}$
@@
5 Experiments
In this section, we evaluate the performance of the proposed algorithms through a series of simulated experiments. The experiments contain two parts, the first of which focuses on the numerical results of synthetic cases while the second is conducted in the Airsim simulator [40] to test the performance in the realistic virtual environment.
5.1 Experiments on synthetic cases
@@
5.1.1 Brute force experiments
@@
5.1.2 Different range and angle experiments
@@
5.2 Experiments in the simulator
In addition to the tests on synthetic data, we carried out our algorithms in the Airsim simulator for practical applications, which is an open-source simulator developed by Microsoft for training and testing autonomous vehicles and drones in realistic virtual environments. The simulation consists of two parts. Firstly, We create 50 cylinders as objects in the 3DsMax software with radius of 1 m and height of 2 m. Subsequently, we paste on the cylinders high definition English advertisement images downloaded from the Internet, which are then imported into the simulator (see Figure 4). The map size is set to 150 m and 30 cases are generated for each $n$ in $\left\{5,10,15,20\right\}$. For each case, $n$ objects are randomly chosen from the 50 ones and $q^{*}$ is set to $\left\{0.3,0.5,0.7,0.9\right\}$. The drone will fly along the paths produced by different order determination algorithms and dynamic programming algorithm. Additionally, the initial paths are included provided by RS, NPF, and GTSP without adjustment. We also add a path that yields the highest observation quality from all the objects. The path is solved by using the algorithm from [32] on the observation points that produces maximum observation quality of each object. During the flight, the UAV takes photos of the objects which are sent to the CnOCR API in Python [41]. The observation quality is reflected in the accuracy of the recognized words in the photos.
(a) ’WORLDSFORESTS’
13 / 13
(b) ’AdveStatis’
10 / 21
(c) ’DLENDHAND’
9 / 21
Furthermore, a virtual city environment was constructed by using the ’Stylized Town Package’ and ’Billboards VOL.1’ packages from Unreal Engine Market [42]. Dozens of different billboards and buildings with boards of names were randomly placed in the map to simulate a real and complex city (see Figure 5). The performance of the algorithms were then validated in this environment by setting $n$ to 15 and treating the boards of the buildings and billboards as objects. The flight height is slightly higher than the tallest building to avoid collision. We ran 15 cases in total, with the concrete processes and goals same as the first part.
Figure 5: Virtual city environment in the simulator
To calculate the recognition accuracy in one case, we solve the longest common subsequence between the recognition result and the label of the object as the correctly recognized string. Suppose that the labels of the objects in one case are =$\mathcal{L}=\left\{l_{1},l_{2},...,l_{m}\right\}$, the correctly recognized strings $\mathcal{C}=\left\{c_{1},c_{2},...,c_{m}\right\}$. We define the accuracy as the sum of all correctly recognized letters over all the letters in the labels, $\frac{\sum_{i}{|c_{i}|}}{\sum_{i}{|l_{i}|}}$. Figure 4 shows the photos taken during the flight and the according OCR results.
Experiments on synthetic cases consider that the observation distance and angle are equal for all objects. In the simulation trails, however, owing to the different size of the images and words, the observing distance and observing angle are calculated separately for each object. We stipulate that the UAV should recognize at least $30\%$ correct letters for each object from some observation point. Based on this, $d_{max}$ and $d_{min}$ are searched via a binary search algorithm, after placing the UAV right on the front of the object. The observing angle $\theta$ is then obtained from a second binary search algorithm that guides the UAV to fly on the arc with the radius of $d_{max}$. Figure 6 is an overview perspective showing the different observation range of objects and Table 1 contains a proportion of the objects’ information.
@@ 5.2 Experiments in the simulator
Table 1: Labels, maximum and minimum observation distance and maximum observation angles of some objects
@@
5.3 Numerical results
In this section, the statistic results are presented in the following figures and tables. We first present the results of the brute force experiments, which are the comparison of our order determination algorithms with the optimal order. The results of different range and angle experiments are then demonstrated before presenting the running time. Finally, we show the simulation results.
5.3.1 Comparison with brute force
@@
5.3.2 Different $n$ and range
We fix the observation range to 4, 8 and 12, and plot Figure 7, in which the effect of $n$ is illustrated. After that, $n$ is fixed and the effect of observation range is shown in Figure 8.
(a) observation range = 4
(b) observation range = 8
(c) observation range = 12
@@ 5.3.2 Different $n$ and range
(a) object number = 10
(b) object number = 20
(c) object number = 30
@@
When $d_{max}=4$, the overall approximation ratio has a declining tendency with 1.61 when $n$ is 5 and 1.31 when $n$ is 30 for the GTSP method. Other order determination methods produce similar results, and the relative superiority is shown to be the same as that in the brute force section. NPF achieves 1.71 approximation ratio when $n$ is 5 which gradually declines to 1.58 when $n$ is 30. The results bear some differences for larger observation ranges. When the range is 8 and 12, the approximation ratios both undergo a decline when $n$ is smaller than 15 and then experience a rebound. Concretely, for cases where the range equals 8, The GTSP methods achieves the approximation ratio of 1.69 when $n$ is 5, which gradually descends to 1.45 when $n$ is 15. The ratio then rebounds to 1.49. When $d_{max}$ is 12, the approximation ratio ranges from 1.51 to 1.98, slightly higher than the overall situation in $d_{max}=8$. The performance of NPF method seems to be susceptible to the growing $n$ but the other method remains relatively stable. In another word, our algorithms are robust to the growing number of objects.
We then analyse the effect of observation range. The number of objects is selected to be respectively 10, 20, and 30. It is apparent to notice the growing trend in approximation ratio with the rise in the observation range from 4 to 10 for all $n$. The ratio then remains stable when the range further increases to 12. When $n$ is 10, GTSP method achieves the approximation ratio ranging from 1.41 to 1.59, roughly 0.1 or 0.2 smaller than the other methods. The results bear a resemblance for large $n$, in which the ratio increases from 1.32 to 1.52 when $n$ is 20 and 1.32 to 1.6 when $n$ is 30. As a brief summary, there is a modest increase in approximation ratio along with the observation range and number of objects, in that the algorithms probably have slight difficulty in solving complex cases.
5.3.3 Effectiveness of dynamic programming algorithm
According to the experimental results of the previous two sections, it can be seen that the dynamic programming algorithm can achieve optimal performance combined with the GTSP observation order. This part of the experiment compares our algorithm with 4 baseline algorithms. The baseline methods are: Random Selection (RS), Nearest Point First (NPF) [43], Generalized Traveling Salesman Problem (GTSP) [39], and Maximum Observation Quality Path (MaxQ). Since the TSPO and LBTSP methods only generate the order of observed objects rather than a feasible path, they are not used as baseline algorithms for comparison in this part.
@@ 5.3.3 Effectiveness of dynamic programming algorithm
We have counted the percentage of cases in which RS, NPF, and GTSP methods can achieve the quality requirement without the assistance from our dynamic programming algorithm, the effectiveness and the necessity of which is to be explored. TSPO and LBTSP methods are omitted since they only yield the visiting order instead of feasible paths. For those cases in which the order determination methods have already been capable of achieving the quality requirement, we calculate the percentage by which the dynamic programming algorithm is able to reduce the path lengths.
@@
The number of objects is selected to be 5, 15, and 25. As the quality requirement grows, the general tendency is extremely in accordance with the expectation. Concretely, the satisfying percent of all the three methods remains high when $q^{*}\leq 0.5$, with GTSP order performing slightly poorly than the others. For stricter $q^{*}$, however, the percentage drops rapidly to $20\%$ and finally to 0, which is given by NPF method when $q^{*}$ is 0.9. Based on the outcomes above that the order determination algorithms can barely find satisfying paths for large $q^{*}$, it is necessary to use dynamic programming to improve the path in order to fulfill the quality requirement.
As shown in Figure 9, our algorithm solves a shorter path than the baselines. Since the baseline algorithms did not consider observation quality, the paths are presented as straight lines in the figure. Compared to other algorithms, the algorithm proposed in this study can obtain significantly shorter paths. For example, when $n$ is 5, the average path length of the baseline algorithm RS is about 521m, NPF is 524m, and GTSP is about 509m, while the average path length of the algorithm proposed in this study is only 491m, and it can also meet the perceptual quality constraint. When seeking relatively higher perceptual quality, there is a trade-off between perceptual quality and path length. Compared to the MaxQ path in the figure, the algorithm in this study can search for a path with $90\%$ of the maximum perceptual quality but with a significantly shorter path length, reducing the original path by about 79m.
(a) object number = 5
(b) object number = 15
(c) object number = 25
Figure 9: Path length of algorithms under different observation quality constraints
5.3.4 Simulation results